# An Adaptive Gradient Regularization Method

### v3

@@ 4. AGR Properties
In the following experimental results, we will see how AGR embedded into Adan or AdamW improves the generalization performance and training speed across multiple neural networks. In this section, let us give a detailed theoretical analysis of why it works well.
@@
4.1 Adaptive Gradient Regularization
Smooth the loss landscape It has been proved that batch normalization(BN) (Ioffe and Szegedy, 2015) and weight standardization(WS) (Qiao et al., 2019) can smooth the loss landscape during the training process by reducing the Lipschitz constants of the loss and its gradients. GC (Yong et al., 2020a ) also performs a theory analysis that this method can make the gradient of weight and the Hessian matrix of weight more stable. Specifically, GC (Yong et al., 2020a ) use the magnitude of $\left|\left|\nabla_{\textbf{w}}\mathcal{L}\right|\right|_{2}$ and $\left|\left|\nabla_{\textbf{w}}^{2}\mathcal{L}\right|\right|_{2}$to capture the Lipschitzness of loss and its gradient, respectively. BN and WS can smooth the loss landscape with operation on weight, providing better Lipschitz properties by indirectly constraining the loss function and its gradients. GC smooths the loss landscape by direct operation on gradients. We also use this mathematical property to give a theory analysis of the AGR method as below and obtain a similar conclusion.
@@ 4.1 Adaptive Gradient Regularization
Theorem 4.1 $\nabla_{\textbf{w}}\mathcal{L}$ is the gradient of loss function $\mathcal{L}$ w.r.t. weight w, with the function $\Psi(\nabla_{\textbf{w}}\mathcal{L})$ defined in Eq.(2), we can give the following conclusion about the loss function and its gradient:
$\left\{\begin{array}[]{lr}\left|\left|\Psi(\nabla_{\textbf{w}}\mathcal{L})\right|\right|_{2}\leq\left|\left|\nabla_{\textbf{w}}\mathcal{L}\right|\right|_{2},&\\ \left|\left|\nabla_{\textbf{w}}\Psi(\nabla_{\textbf{w}}\mathcal{L})\right|\right|_{2}\leq\left|\left|\nabla_{\textbf{w}}^{2}\mathcal{L}\right|\right|_{2},&\\ \end{array}\right.$ (4)
The proof of theorem 4.1 is attached in the Appendix. It shows that the loss $\mathcal{L}$ and its gradient $\nabla_{\textbf{w}}\mathcal{L}$ are restrained with the AGR method. Like BN, WS, and GC, our method provides better Lipschitzness by implementing a dynamic regularization on the gradients.
4.2 Adaptive Learning Rate
@@
Adjust learning rate depending on gradient During the weight iteration process, the learning rate and gradient contribute to the update direction and magnitude together (Cutkosky et al., 2023). The first-order momentum and second-order momentum are introduced to accelerate loss descending and scaling learning rate, respectively. In addition, the adaptive learning rate depends on the gradient magnitude. In other words, the learning rate in each iteration for a weight is integrated with its gradient. From this perspective, our AGR method also integrates an adaptive learning rate with a specific gradient.
Theorem 4.2 AGR can adapt learning rate $\eta_{i,j}$ integrated with the gradient $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$, their adaption tendency keep consistent.
The proof of theorem 4.2 is attached in the Appendix. From this perspective, we can understand that the operation on gradients and the learning rate are not separate. It is a better way to view why the AGR works well. Hence, we can expect faster training speed and better generalization performance due to the smoother loss landscape.
5.    Experimental Result
We apply our AGR method to vision tasks and natural language processing (NLP) tasks. Vision tasks consist of image generation and classification. In the image generation task, we evaluate the AGR method to the representative SOTA backbone U-Net (Ronneberger et al., 2015) in the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). In the image classification task, we mainly investigate the ARG method under the conventional supervised setting, including CNN-type architectures: ResNets (He et al., 2016), VGG (Iglovikov and Shvets, 2018), ConvNext (Liu et al., 2022) and transformer-type architectures: ViT (Yin et al., 2022; Yuan et al., 2021) and Swin (Liu et al., 2021; Liang et al., 2021). In the NLP task, we evaluate the AGR method in ABERT. (Devlin et al., 2018; Lan et al., 2019)
We focus on applying ARG to the SoTA optimizer in the architectures mentioned above to evaluate their performance. For example, In the image classification task, the default/SoTA optimizer is AdamW (Loshchilov and Hutter, 2017) in CNN-type architectures and ViTs. In the NLP task, the default/SoTA optimizer is LAMB and we set the optimizer as AdamW. Except for that, we also apply the AGR method to Adan (Xie et al., 2022) optimizer, which achieves better performance than AdamW in U-Net architecture under the image generation task. Restricted to the computational resource, we do not implement the ARG modification in all optimizers like SGD and Adam under all tasks. We also focus on AdamW because SGD and Adam achieve much worse accuracy than AdamW in image classification and generation tasks.
@@ 5.1 Experimental Setup
For evaluating and validating the AGR method generally and comprehensively, we conduct our experiments across image generation, image classification, and NLP tasks across multiple popular neural networks and open-source datasets. We arrange our experiments as follows:
  • We first conduct an image generation task on the CIFAR10 dataset (Krizhevsky et al., 2010), which consists of 10 classes, and each class has 5000 images for training and 1000 for testing. The image resolution is 32 $\times$ 32. In this task, we use the famous diffusion model paradigm: Denoising Diffusion Probabilistic Model (DDPM) with U-Net architecture as a solver for training and apply the AGR to the Adan optimizer. For comparison, we also conduct experiments with other popular optimizers.
  • Next, we conduct a traditional image classification task on the CIFAR100 dataset (Lin et al., 2013), and Tiny-ImageNet (Le and Yang, 2015) across multiple classical neural networks (VGG, ResNet, ViTs, ConvNext, etc.). Due to the restricted resources, we only train the tiny version of ViTs and ConvNext. The CIFAR100 dataset consists of 100 classes, each with 500 images for training and 100 for testing. The image resolution is 32 $\times$ 32. The Tiny-ImageNet consists of 200 classes, each with 500 images for training, 50 for validating, and 50 for testing. The image resolution is 64 $\times$ 64. Due to the images in the test dataset not having class labels, we use the validation dataset for evaluation.
  • Finally, we conduct language representation tasks on the WikiText-2 dataset (Radford et al., 2019). Due to the restricted resources, we train Albert (Lan et al., 2019), a lite version of Bert, for masked language model (MLM) and sentence order prediction (SOP) tasks.
  • For better evaluation, each experiment is conducted four times, and then we calculate their mean as the final result.
If there is no additional note, we apply the AGR method to the fully connected layers and convolutional layers across all neural architectures in this work. For Adan and AdamW, we keep all default sets while using the LambdaLR learning rate scheduler if comparison experiments are not conducted on the hyperparameters. There is no integrated and additional operation on the optimizers except for the simple two codes representing the AGR method. All experiments are conducted on NVIDIA A40 with Pytorch framework as a 2.0 version.
@@ 5.2 Generative Model: DDPM
In this section, we present the results of DDPM training on the CIFAR10 dataset with different optimizers. We evaluate the performance with FID and IS scores, respectively., we also calculate the standard deviation of the IS score of generating 50000 images. For each experiment, we train DDPM with 2000 epochs and a batch size of 128. To save time, we introduce the denoising diffusion implicit model (DDIM) (Song et al., 2020) with steps as 100 for acceleration generation in each evaluation. As shown in Table 1 and Table 2, the Adan optimizer achieves the best performance among the popular optimizers, which also certifies its efficiency in the image generation task. Notably, Adan(AGR) improves performance with a lower FID score and higher IS score. It shows that the AGR method improves the performance of U-Net architecture in predicting noise, which verifies that AGR is an efficient optimizer development technique for image generation. We also consider transferring this method to other diffusion model paradigms in future work to verify its more general performance. We need to stress that we focus on the comparison of performance after embedding the AGR method under the same setting. Hence, we don’t pay attention to the hyperparameter tuning for more accuracy.
@@ 5.3 Supervised Classification on TinyImageNet and CIFAR100
In this section, we conduct the traditional image classification task on the CIFAR100 and Tiny-ImageNet datasets. We train VGG11, ResNet-18, ResNet-50, and ResNet-101 architectures on the CIFAR100 dataset and train ConvNext and ViTs architectures on the Tiny-ImageNet dataset. We focus on applying the AGR on the AdamW optimizer. On the CIFAR100 dataset, each experiment is conducted in 200 epochs on 2 GPUs with batch size 64 per GPU. On the Tiny-ImageNet dataset, each experiment is conducted in 300 epochs on 2GPUs with batch size 128 per GPU.
@@ 5.3.1 Results on CIFAR100
In this section, we evaluate the AGR method with some classical CNN- architectures on the CIFAR100 dataset, and we can see that the Top-1 accuracy can be improved to different degrees. As Figure 2 shows, the training loss dives deeper with the AGR method, and the test accuracy is improved and more smooth, which shows that the AGR method can stabilize the training process and improve the general performance in the ResNet18 backbone. We also apply the AGR to AdamW on other popular backbones such as ResNet50 and ResNet101. As Figure 3 shows, the training loss landscape of VGG11 with the AGR method is smoother with higher performance. Hence, it achieves better accuracy in the test datasets in a smoother way, which verifies that the AGR can accelerate the training process and improve the generalization ability of AdamW in the VGG11 backbone. We also present the evaluation result of the AGR on the ResNet18 backbone. As Table 3 shows, the AGR method could improve the performance of around 1% of AdamW on them.
@@ 5.3.2 Results on Tiny-ImageNet
In this section, we evaluate the AGR method on transformer architectures and ConvNext on the Tiny-imageNet dataset. For a more comprehensive comparison, we apply the AGR to multiple variants of TinyViT, Swin, and ConvNext. Specifically, there are three variants of TinyVit with 5M, 11M, and 21M parameters. There are also tiny versions of ConvNext and Swin, respectively. As Figure 3 shows, the AGR method could reduce the training loss and improve the test accuracy mostly through the entire training process. The test accuracy of Swin-Tiny is improved more in the later epochs of training, which also improves the generalization capability and accelerates the training process. As Table 3 shows, the AGR method can improve the performance >0.5% in general on ConvNext and Swin backbones. As Table 4 shows, the test accuracy in TinyViT variants is improved by 1.7% at most. Similar to the multiple ResNet variants, TinyViT with more parameters cannot also excavate more pattern information due to the limited training images which achieves a little higher accuracy than TinyViT with fewer parameters. Hence, the AGR method can be embedded into multiple neural networks in the transformer architecture.
@@ 5.4 Language Representations: ALBERT
In this section, we evaluate the AGR method on the natural language representation tasks with the lite version of BERT: Abert. The language representation tasks consist of masked language model (MLM) and sentence order prediction (SOP). We evaluate the performance of the WikiText-2 dataset on a single GPU. Each experiment is conducted in 25 epochs with batch size 16. In this section, we also evaluate the AGR method across different weight decay in AdamW. As Table 6 shows, the AGR method can improve the SOP accuracy across different weight decay, and we also see the importance of weight decay in the model training process.
5.5 Ablation Study
Through extensive experiments, we found that the AGR method cannot improve the training performance during the entire process. According to our theoretical analysis, the gradient and learning rate will be reduced adaptively based on the current gradient magnitude. The learning rate will decrease more when the gradient magnitude is larger. However, the learning rate needs to be larger in the later training epochs. Hence, the performance will be better when suspending the AGR method after a period of training. In this section, we conduct experiments on image classification tasks with the TinyViT-11M backbone and evaluate the influence of the AGR running epochs. As Table 7 shows, when running the AGR method with 250 epochs instead of entire epochs, we can achieve higher accuracy. This verifies our analysis that the learning rate in the later training epochs should be kept larger.
@@ 6. Conclusions
We proposed a new optimization technique that normalizes the gradient vector as its coefficient vector based on the gradient magnitude and subtracts the dot of the gradient vector and its coefficient vector from the vanilla gradient vector. Hence, the gradient can be regularized in an adaptive way named adaptive gradient regularization (AGR). From theory analysis, we show that the AGR can improve the Lipschitzness of the loss function. Then, extensive experiments on image generation, image classification, and natural language processing show that the AGR can stabilize the training process and accelerate the training with better generalization.
Appendix A Appendix
A1. Theorem 4.1 Proof
Proof: For weight $w_{i,j}$ in weight vectors and definition in eq(1) and eq(2), we have:
$\begin{split}\left|\left|\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})\right|\right|^{2}_{2}=&\left|\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}\leq\left|\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}\\ \end{split}$ (5)
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})=&\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})\\ \end{split}$ (6)
Next, we need to discuss about two cases with $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$<0 and $\nabla_{\textbf{w}_{i,j}}\mathcal{L}\geq 0$
when $\nabla_{\textbf{w}_{i,j}}\mathcal{L}\geq 0$:
$\begin{split}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})&=-\nabla_{\textbf{w}_{i,j}}\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}\\ &=-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ \end{split}$ (7)
Combined with eq(6), we have:
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})&=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}-2\nabla_{\textbf{w}_{i,j}\mathcal{L}}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+(\nabla{\textbf{w}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{W}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &\leq\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\end{split}$ (8)
when $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$<0:
$\begin{split}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})&=\nabla_{\textbf{w}_{i,j}}\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}\\ &=\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ \end{split}$ (9)
Combined with eq(6), we have:
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})&=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1+\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}+2\nabla_{\textbf{w}_{i,j}}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+(\nabla{\textbf{w}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{W}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &\leq\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\end{split}$ (10)
In the optimization process, we need to find the optimal local minimum as the global minimum. Hence, the landscape of loss containing these local minima is convex. So we have:
$\begin{split}\left|\left|\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})\right|\right|^{2}_{2}\leq\left|\left|\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}\end{split}$ (11)
A2. Theorem 4.2 Proof
In this proof, we take optimizer AdamW with AGR to illustrate theorem 4.2 is certified. From the definition of momentum $m_{t}$ in AdamW at t-st update for a specific weight, we can derive recursively:
$m_{t}=\sum^{t-i}_{i=0}\beta^{i}_{1}(1-\beta_{1})(\Psi(\nabla_{\textbf{w}_{t}}\mathcal{L}))_{t-i}$ (12)
From the simplicity we don’t apply ARG to the second-order momentum, we have the following weight update process:
$\begin{split}w_{t+1}&=w_{t}-\eta_{t}m_{t}\\ &=w_{t}-\eta_{t}\sum^{t-i}_{i=0}\beta^{i}_{1}(1-\beta_{1})(\Psi(\nabla_{\textbf{w}_{t}}\mathcal{L}))_{t-i}\\ &=w_{t}-\eta_{t}\sum^{t-i}_{i=0}\beta^{i}_{1}(1-\beta_{1})((1-\alpha_{t-i})(\nabla_{\textbf{w}_{t}}\mathcal{L}))_{t-i}\\ \end{split}$ (13)
Assume for each time $\alpha_{t-i}=\hat{\alpha}$ is the same for clearer certification, we have:
$\begin{split}w_{t+1}&=w_{t}-\eta_{t}\sqrt{(1-\hat{\alpha}^{t})}\sum^{t-i}_{i=0}\beta^{i}_{1}(1-\beta_{1})\sqrt{(1-\hat{\alpha}^{t})}(\nabla_{\textbf{w}_{t}}\mathcal{L})_{t-i}\\ \end{split}$ (14)
Through this perspective, we can see the same tendency in learning rate and gradient adaption with AGR.

### v4

@@ 4. AGR Properties
In this section, we shall provide a theoretical analysis to explain why the studied adaptive gradient regularization can accelerate the training efficiency and improve the generalization performance of deep neural networks in different tasks.
@@
Smoothing the loss landscape: The loss landscape of deep neural network provides a valuable perspective for investigating the training stability, robustness and generalization of neural networks. Note that we can evaluate the loss landscape by the Lipschitzness and gradient of the loss function. In particular, we use the magnitude of $\left|\left|\nabla_{\textbf{w}}\mathcal{L}\right|\right|_{2}$ and $\left|\left|\nabla_{\textbf{w}}^{2}\mathcal{L}\right|\right|_{2}$to capture the Lipschitzness of loss and its gradient, respectively. We also note that a better Lipschitzness of the loss with smaller $\left|\left|\nabla_{\textbf{w}}\mathcal{L}\right|\right|_{2}$, often indicates a smoother landscape and more effective training performance.
@@
Theorem 4.1 $\nabla_{\textbf{w}}\mathcal{L}$ is the gradient of loss function $\mathcal{L}$ w.r.t. weight w, with the function $\Psi(\nabla_{\textbf{w}}\mathcal{L})$ defined in Eq.(1). We have the following results about the loss function and its gradient:
$\left\{\begin{array}[]{lr}\left|\left|\Psi(\nabla_{\textbf{w}}\mathcal{L})\right|\right|_{2}\leq\left|\left|\nabla_{\textbf{w}}\mathcal{L}\right|\right|_{2},&\\ \left|\left|\nabla_{\textbf{w}}\Psi(\nabla_{\textbf{w}}\mathcal{L})\right|\right|_{2}\leq\left|\left|\nabla_{\textbf{w}}^{2}\mathcal{L}\right|\right|_{2}.&\\ \end{array}\right.$ (4)
The proof of theorem 4.1 is attached in the Appendix. Intuitively, the loss $\mathcal{L}$ and its gradient $\nabla_{\textbf{w}}\mathcal{L}$ using AGR, can be restricted by a tighter bound. Hence, we naturally can improve the Lipschitzness of the loss function and its gradient in training deep neural networks, by dynamically regularizing gradients with AGR.
@@
Adjusting learning rate using gradients: During the update of weight, the learning rate and gradient contribute to the update direction and magnitude together (Cutkosky, Defazio, and Mehta 2023). The first-order momentum and second-order momentum are introduced to accelerate loss descending, and to rescale the learning rate, respectively. In addition, the learning rate can also be affected by the gradient magnitude. In other words, the learning rate for a weight at each iteration, can be adaptively adjusted according to its gradient, using the AGR method.
Theorem 4.2 AGR adaptively adjusts the learning rate $\eta_{i,j}$ using the gradient $\nabla_{\textbf{w}_{i,j}}\mathcal{L}$.
The proof of theorem 4.2 is attached in the Appendix. Note that AGR can simutaneously regularize the gradients and effectively adjust the learning rate, via Eq.(2). Intuitively, AGR regularizes the gradient when it is too large, and thus naturally decreases the learning rate, accordingly. We thus expect to achieve faster training speed, and better generalization performance by obtaining the smoother loss landscape using the AGR method.
5.    Experimental Results
We evaluate the effectiveness of the AGR method in improving the training efficiency and generalization performance for neural networks, using various tasks including image classification, image generation, and language representation. For the image generation, we applied the AGR method to speed up the training of the U-Net architecture (Ronneberger, Fischer, and Brox 2015) of the denoising diffusion probabilistic model (DDPM) (Ho, Jain, and Abbeel 2020). For the image classification, we mainly evaluated the AGR method under the conventional supervised setting, e.g., in speeding up the training of CNN architectures: ResNets (He et al. 2016b), VGG (Iglovikov and Shvets 2018), ConvNext (Liu et al. 2022) and Transformer architectures: ViT (Yin et al. 2022; Yuan et al. 2021) and Swin (Liu et al. 2021; Liang et al. 2021). For the language representation task, we applied the AGR method in speeding up the training of ALBERT (Devlin et al. 2018; Lan et al. 2019). More specifically, we evaluate the effectiveness of the AGR in improving the training efficiency and generalization performance of neural networks, by applying the AGR to the state-of-the-art optimizers, and comparing the improvements brought by the AGR. For example, In the image classification task, the default/SOTA optimizer is AdamW (Loshchilov and Hutter 2017) in CNN-type architectures and ViTs. For the NLP task, the default/SOTA optimizer is LAMB, and we set the optimizer as AdamW. In addition, we also applied the AGR method to Adan optimizer (Xie et al. 2022), which achieves better performance than AdamW in training U-Net architecture for the image generation. We considered to apply AGR to AdamW, and to compare it with AdamW instead of Adam and SGD, because SGD and Adam achieve much worse performance than AdamW, for image classification and generation.
@@ 5.1 Experimental Setup
To thoroughly evaluate the AGR method, we considered to conduct the experiments in training various neural architectures, using multiple representative datasets.
More specifically, the experimental setup are detailed as follows:
  • We conducted image generation task using the CIFAR10 dataset (Krizhevsky, Hinton et al. 2010), which consists of 10 classes, and each class has 5000 images for training and 1000 for testing. The image resolution is 32 $\times$ 32. In this task, we used the famous diffusion model paradigm: Denoising Diffusion Probabilistic Model (DDPM) with U-Net architecture as a solver for training and applied the AGR to the Adan optimizer. Moreover, we also included the other popular optimizers in the comparison.
  • We conducted image classification using multiple classical neural networks (VGG, ResNet, ViTs, ConvNext, etc.), on the CIFAR100 dataset (Lin, Chen, and Yan 2013), and Tiny-ImageNet (Le and Yang 2015). For the experiments, we trained the tiny version of ViTs and ConvNext. The CIFAR100 dataset consists of 100 classes, each of which has 500 images for training and 100 for testing. The image resolution is 32 $\times$ 32. The Tiny-ImageNet consists of 200 classes, each of which has 500 images for training, 50 for validating, and 50 for testing. The image resolution is 64 $\times$ 64. As the labels of the images are not provided in the test set, we thus use the validation dataset for evaluation.
  • Finally, we conducted language representation tasks on the WikiText-2 dataset (Radford et al. 2019). For the experiments, we trained ALBERT (Lan et al. 2019), a light version of BERT, for masked language model (MLM) and sentence order prediction (SOP) tasks.
  • For the better evaluation, each of the experiments is conducted four times, and then we calculated the mean to report the final results.
Unless otherwise specified, we applied the AGR method to the fully connected layers and convolutional layers of the neural architectures in the experiments. For Adan and AdamW, we used their default settings while using the LambdaLR learning rate scheduler if comparison experiments are not conducted on the hyperparameters. There is no integrated and additional operation on the optimizers except for the simple three lines of codes representing the AGR method. All the experiments are conducted on NVIDIA A40 with Pytorch framework(2.0 version).
@@ 5.2 Generative Model: DDPM
In this section, we present the results of DDPM training on the CIFAR10 dataset using the compared optimizers. We evaluate the performance by FID and IS scores. We also calculated the standard deviation of the IS score in generating $50,000$ images. For each of the experiments, we trained DDPM with $2,000$ epochs with a batch size being $128$. To save time, we introduce the denoising diffusion implicit model (DDIM) (Song, Meng, and Ermon 2020) with steps being set to $100$, to accelerate the generation in each evaluation. As shown in Table 1 and Table 2, the Adan(AGR) optimizer achieves the best performance among the compared optimizers, which also validates its effectiveness in the image generation task. It means that the AGR method can improve the learning performance of U-Net architecture, which validates that AGR is an effective optimization technique for image generation. We also consider to employ this method in the training of other diffusion models in future work, to validate its effectiveness. Note that we only consider the comparisons using the AGR method under the same setting. Hence, we omit the hyperparameter tuning, which may lead to achieving better accuracy.
@@ 5.3 Supervised Classification on TinyImageNet and CIFAR100
For the image classficaiton tasks, we applied the AGR to the AdamW optimizer, and evaluated the improvements bround by AGR. For the CIFAR100 dataset, each of the experiments is conducted with $200$ epochs on $2$ GPUs with batch size being $64$ per GPU. For the Tiny-ImageNet dataset, each of the experiments is conducted with $300$ epochs on $2$ GPUs. The batch size is $128$ per GPU.
@@ 5.3.1 Results on CIFAR100
We evaluated the AGR method in the training of some classical CNN- architectures on the CIFAR100 dataset. We can see that how much the improvments brought by AdamW(AGR) relative to AdamW is, for the Top-1 accurarcy. As shown in Figure 2, the training loss goes deeper with the AGR method, and the test accuracy is improved and smoother, which shows that the AGR method can stabilize the training and improve the learning performance for the ResNet18 backbone. We also applied the AGR to AdamW for other popular backbones such as ResNet50 and ResNet101. As shown in Figure 3, the training loss of VGG11 is lower, and test accuracy is higher, using the AGR method. Hence, it achieves better accuracy in the test datasets in a smoother way, which validates that the AGR can accelerate the training, and improve the generalization performance of AdamW, for the VGG11 backbone. We also present the result of the AGR for the ResNet18 backbone. As shown in Table 3, the AGR method can improve the performance of AdamW with around $1\%$ improvements for the Top-1 accuracy.
@@
Table 8: FID score of DDPM on CIFAR10
Epochs 0 400 800 1200 1600 2000
Adan(AGR) 7.73 7.64 7.58 7.02 7.81 7.44
@@ 5.3.2 Results on Tiny-ImageNet
We conducted the experiments on the Tiny-ImageNet. To evaluate the efficiency and generalization performance of the AGR, we applied the AGR method to train Transformer architectures and ConvNext using the Tiny-imageNet dataset. For a more comprehensive comparison, we applied the AGR to the multiple variants of TinyViT, Swin, and ConvNext. Specifically, there are three variants of TinyVit with 5M, 11M, and 21M parameters. As shown in Figure 3, the AGR method can further decrease the training loss, and improves the test accuracy, through the entire training. The test accuracy of Swin-Tiny is improved with a large margin in the later epochs. As shown in Table 3, the AGR method can improve the performance with >$0.5\%$ improvement, for the ConvNext and Swin backbones. As shown in Table 4, the test accuracy for the TinyViT variants, is improved by 1.7% at most. Similar to the multiple ResNet variants, TinyViT with more parameters cannot also excavate more pattern information due to the limited training images which achieves a little higher accuracy than TinyViT with fewer parameters. Hence, the AGR method can be embedded into multiple neural networks in training the Transformer.
@@ 5.4 Language Representations: ALBERT
Here we applied the AGR method to train the light version of BERT: ALBERT. The language representation tasks consist of masked language model (MLM) and sentence order prediction (SOP). We evaluated the performance of the compared optimizers using the WikiText-2 dataset on a single GPU. Each of the experiments is conducted with 25 epochs. The batch size is 16. In this section, we also considered to evaluate the AGR method, by using multiple weight decays in AdamW. As shown in Table 6, the AGR method can improve the SOP accuracy across different weight decays, and we also see the importance of weight decays in training the neural architectures.
5.5 Ablation Studies
Through the extensive experiments, we found that the AGR method cannot improve the training performance during the entire process. According to our theoretical analysis, the gradient and learning rate will be reduced adaptively based on the current gradient magnitude. The learning rate will decrease more when the gradient magnitude is larger. However, the learning rate needs to be larger in the later training epochs. Hence, the performance will become better when suspending the AGR method after a period of training. Here, we conducted experiments on image classification tasks using the TinyViT-11M backbone, and image generation tasks using the DDPM, to evaluate the influence of the AGR over the epochs. As shown in Table 7, when we ran the AGR method till the 250 epochs instead of the entire epochs, we can achieve higher accuracy. This validates our assumption that the learning rate in the later training epochs should be set to larger, relative to the epochs before. When we ran the AGR method with $1,200$ epochs instead of the entire training, we can obtain the best generalization performance.
@@ 6. Conclusions
This paper proposed a new optimization technique to improve the training efficiency and generalization performance for deep neural networks, in a simple-yet-effective way. In particular, theoretical results are provided to guarantee that the new algorithm can effectively smooth the loss landscape and adaptively adjust the learning rate. We conducted various deep neural network-related experiments to demonstrate the superior performance of the novel optimization algorithm, compared with the state-of-the-art optimization methods. In the near future, we will further explore the effectiveness of this novel optimization technique in accelerating the training efficiency and robustness for Transformer and Mamba, which may significantly benefit the research on large language models and video generations.